Tangent plane at any point of sphere x^2+y^2+z^2=r^2 meets the coordinate axis at A,B, C. Show that the locus of the point of intersection of planes drawn parallel to the coordinate planes through A,B,C is the surface x^-2+y^-2+z^-2=r^-2


1
Expert's answer
2021-03-01T07:16:28-0500

The equation of the tangent plane at any point "(x_0,y_0,z_0)" :

"xx_0+yy_0+zz_0=r^2"

The plane meets the coordinate axes at

"A:(r^2\/x_0,0,0),B:(0,r^2\/y_0,0),C:(0,0,r^2\/z_0)"

The equations of the planes parallel to the coordinate planes through A, B, C:

"x=r^2\/x_0,y=r^2\/y_0,z=r^2\/z_0"

Then:

"x^{-2}+y^{-2}+z^{-2}=\frac{1}{r^2}(x_0^2+y_0^2+z_0^2)=\frac{r^2}{r^4}"

"x^{-2}+y^{-2}+z^{-2}=r^{-2}"